OpenStudy (anonymous):

Help! Solve √7 -3x = 1 every time I do this I get a different answer :(

OpenStudy (anonymous):

transfer sqrt(7) to the other side multiply both sides by (-1) solve for x

OpenStudy (anonymous):

I should clarify the 7-3x is all under the root

OpenStudy (anonymous):

oh sorry. My connection got lost. and I didnt know. so... sqrt(7-3x) = 1 square both sides: (sqrt(7-3x))^2 = 1^2 <-- like so 7-3x = 1 -3x = 1-7 -3x = -6 3x = 6 x = 2

OpenStudy (anonymous):

@northwestbon

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

You did understand why it was squared both sides right? and no probs.

OpenStudy (anonymous):

yes i wound up pounding through it and figured it out. But thank you for at lease proving i did do it right in the end =D

OpenStudy (anonymous):

lol Well atleast you got the answer.